OpenStudy (math2400):

could someone help me with 7 and 11 on this? Natural logs w/derivatives

OpenStudy (anonymous):

where is 7 and 11?

OpenStudy (math2400):

sorry just posted it

OpenStudy (math2400):

for 7 i was able to do this: log x^2-3x = 1 but i'm not sure what to do next

OpenStudy (anonymous):

log (x) + log (x-3) = 1 First step is to combine the left side into ONE logarithm. The left side is just log (x(x-3)) = 1 Now write the log equation in exponential form, you get x(x-3) = 10 x^2 - 3x = 10 x^2 - 3x - 10 = 0 (x-5)(x+2) = 0 x = 5 and x= -2 Reject x = -2 as it doesnt check int the original equation. x = 5 is the only solution.

OpenStudy (anonymous):

Makes sense?

OpenStudy (math2400):

sorry had to reboot. but yes it does! thank you. Would u happen to be able to help me with 11 though?

OpenStudy (anonymous):

I will start you off......... The derivative of ln (x) is 1/x The derivative of ln (u) = i/u du So, the derivative of ln(lnx) is 1/ln x times 1/x = 1/xlnx Now you want the derivative of ln(ln(lnx)))

OpenStudy (anonymous):

you want the derivative of (1/xlnx) Use quotient rule

OpenStudy (math2400):

got it thanks! I'll work it out and make sure u get it right with the answer key. really appreciate it(:

OpenStudy (anonymous):

Welcome.